OpenStudy (anonymous):

Find the equation of a circle in standard form where C(6,-2) and D(-4,4) are endpoints of a diameter.

OpenStudy (anonymous):

(x-1)^2+(y-1)^2=34

OpenStudy (anonymous):

thanks, but how did you get your answer! :)

OpenStudy (anonymous):

first the center point of circle is found as it is located at the middle of two points you know, (x,y)=((6-4)/2,(-2+4)/2)=(1,1). Then the square of radius is found by ((6-1)^2+(-2-1)^2)=34. Then the standard form can be written as, (x-1)^2+(y-1)^2=34

OpenStudy (anonymous):

ohh ok